OpenStudy (jacobbenvenutty):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of one half? Please don't just give me the answer, I would like to know how to do it for future problems.

OpenStudy (anonymous):

Use point-slope form: (y-k) = m (x-h) where (h,k) is your point and m is your slope. To start you off: (y - (-3)) = 1/2 (x - (-2))

OpenStudy (jacobbenvenutty):

Thanks I remember now